# QMDB: Quick Merkle Database

### v2

@@ 2 Background
LVMT[13] proposes a layered storage model to reduce the space and complexity of maintaining authenticated blockchain states. By partitioning the state into multiple segments and using cryptographic accumulators, it compresses less frequently accessed data while preserving verifiability. Proof generation becomes simpler, as intermediate accumulators shorten authentication paths. However, integrating multiple layers increases system complexity and demands careful configuration—suboptimal settings can lead to poor performance. Furthermore, LVMT depends on well-optimized cryptographic primitives.
@@ 7 Acknowledgments
We gratefully acknowledge the invaluable feedback and assistance provided by the many individuals and teams who helped refine our system. In particular, we thank Patrick O’Grady from Commonware for his expertise and guidance throughout the development of QMDB and the writing of this paper. We also extend our gratitude to Yilong Li and Lei Yang from MegaETH, and Ye Zhang from Scroll, for their insightful review of our design and manuscript. Finally, we thank Robert Habermeier and the Thrum team for their support in conducting the NOMT benchmarks.

### v3

@@ 2 Background
LVMT[13] proposes a layered storage model to reduce the space and complexity of maintaining authenticated blockchain states. To achieve constant-time Merkle root updates, it employs vector commitments and introduces a novel data structure to minimize the costly multiplications involved in vector commitment computations. In addition to the append-only Merkle tree, it incorporates an Addressable Merkle Tree (AMT) to support exclusion and latest-value proofs. However, maintaining an additional AMT with vector commitments increases resource overhead. Moreover, LVMT focuses on ADS design rather than providing a fully unified, verifiable database for end-to-end storage.
@@
MoeingADS  [22] is another novel ADS design featuring an append-only Merkle tree and a memory-efficient representation for exclusion and latest-value proofs. Compared to LVMT, MoeingADS incurs even lower resource overhead while supporting the necessary proof features for ADS usability. However, its compaction and exclusion maintenance are performed in batches only after several blocks, impacting system performance and making it less suitable for stateless validation and Zero-Knowledge Proof-based verification. Additionally, MoeingADS lacks support for streaming transaction handling, a fine-grained pipelining system for higher throughput, and a low-memory indexer.
@@ 7 Acknowledgments
We gratefully acknowledge the invaluable feedback and assistance provided by the many individuals and teams who helped refine our system. In particular, we thank Patrick O’Grady from Commonware for his expertise and guidance throughout the development of QMDB and the writing of this paper. We thank the MoeingADS dev team for their review on throuhgput optimization. We also extend our gratitude to Yilong Li and Lei Yang from MegaETH, and Ye Zhang from Scroll, for their insightful review of our design and manuscript. Finally, we thank Robert Habermeier and the Thrum team for their support in conducting the NOMT benchmarks.